# Rollbaccine: Herd Immunity against Storage Rollback Attacks in TEEs (Technical Report)

### v1

@@ 4. Block Device Crash Consistency
Rollbaccine provides rollback resistance by guaranteeing that, following a rollback attack, the disk always recovers to a benign state. A formal definition of benign state is thus necessary. Concretely, we must formalize block device consistency in the presence of crashes: what are the states to which a block device (i.e. a disk) can recover to after a crash? These are the same states that Rollbaccine must recover to post-rollback.
We begin with the assumption that disk read and write operations ($O$) are atomic, which is consistent with prior work and the inherent properties of disks (brown2002oracle, ; sivathanu2003semantically, ; pillai2014all, ). We also assume that writes to block devices are guaranteed to persist across crashes only when the persistence flags REQ_FUA or REQ_PREFLUSH are used (crashmonkey, ; crashMonkeyAndACE, ; vijayThesis, ; fuaPreflush, ).
@@
Read requests to block $b$ by thread $t$ are written $R_{inv,t}(b)$; responses are $R_{res,t}(b,val)$, where $val$ is the value returned. Write requests are $W_{inv,t}(b,val,sync)$, where the value to write (if any) is val and sync is a tag with one of the following values: REQ_FUA, REQ_PREFLUSH, REQ_FUA|REQ_PREFLUSH, or $\varnothing$. $W_{res,t}(b)$ is the matching response. We assume that blocks are always written to before they are read from.
To define block device crash consistency, we take as our starting point the definitions of Izraelevitz et al (durableLinearizability, ). We will build up to a definition of linearizability before extending it to crashes. We start by defining a sequential history.
@@
This allows us to reason about multi-threaded histories by comparing each thread’s execution to a sequential history. [^1] For programs that issue concurrent operations per thread using async I/O, we can map each physical thread to multiple abstract threads. When multiple threads operate over the same block, we uses the happens-before relationship to order writes and reads on different threads, as this is necessary to determine whether a history satisfies reads-see-writes.
@@
(3) $V_{1}=O_{res}(b)$ and $V_{2}=W_{inv}(b^{\prime},val,sync)$ where $sync$ contains REQ_PREFLUSH, or
@@
Criteria 1 and 4 are standard (linearizability, ); Criterion 2 was introduced for crash-consistent NVMs (durableLinearizability, ) and also applies to crash-consistent block devices. It states that crashes are global events; all events either happen-before or after a crash. Criterion 3 is new and captures the block-and-thread-agnostic semantics of REQ_PREFLUSH: once a REQ_PREFLUSH is invoked, it is only returned by disk once all previous writes from every thread are flushed and persisted, regardless of which blocks they wrote to.
The happens-before relationship allows us to define what each read returns when other threads write to the same block. Each read of block $b$ must return the value of the latest completed write to that same block $b$, as long as there are no crashes in-between (during which writes may be lost). We formalize this as the reads-see-writes property, which is only defined for crash-free periods of history (we call these eras $\mathcal{E}$).
@@
A history $\mathcal{H}$ respects reads-see-writes if $\forall R_{res}(b,val)\in\mathcal{H}$, there is a preceding write invocation $W_{inv}(b,val,sync)$ with that same $val$ such that $\mathcal{H}=\mathcal{H}_{0}W_{inv}\mathcal{E}_{0}W_{res}\mathcal{E}_{1}R_{inv}\mathcal{E}_{2}R_{res}\mathcal{H}_{1}$, and there does not exist another $W_{inv}(b)$ in the eras $\mathcal{E}_{0},\mathcal{E}_{1},\mathcal{E}_{2}$.
Finally, we consider pending invocations: invocations without a matching response. Pending writes in particular require care as they may (or may not) have been processed by the underlying block device and reflected in the next read. We write $(\mathcal{H})$ to be the set of histories generated from $\mathcal{H}$ by inserting matching responses after some pending invocations. This models situations where pending operations have been persisted to disk. In contrast, let $(\mathcal{H})$ be the history generated from $\mathcal{H}$ by removing all pending invocations. This reflects histories where the operation was not persisted.
@@
1) $\mathcal{S}$ respects reads-see-writes
2) $\forall t$, $\mathcal{H}^{\prime}[t]=\mathcal{S}[t]$ (i.e. $\mathcal{H}^{\prime}$ and $\mathcal{S}$ are equivalent)
3) $V_{1}\prec V_{2}$ in $\mathcal{H}^{\prime}$ implies $V_{1}\prec V_{2}$ in $\mathcal{S}$.
@@
(1) if $\mathcal{H}^{\prime}$ contains $W_{inv}(b,val,sync)$ and its matching response $W_{res}(b)$ where $sync$ contains REQ_FUA or REQ_PREFLUSH, then $\mathcal{D}$ must contain $W_{res}(b)$,
@@
The durable cut is a cut of history that contains (1) all writes tagged with persistence flags and (2) any writes that happen-before a write already in the cut. This is where the extra criteria for REQ_PREFLUSH in the happens-before relationship becomes relevant; if a REQ_PREFLUSH completed, then it must be in the durable cut, and any operations that happened-before it must be in the cut as well. The REQ_FUA flag instead simply guarantees that, if an operation completes with that flag, it will necessarily be in the cut.
@@
Assuming only benign system crashes and no random disk corruption, block device crash consistency precisely captures the set of histories produced by a disk (dmIntegrityMailingList, ). We prove that all histories produced by Rollbaccine are block device crash consistent in Appendix B. In this way, Rollbaccine ensures that the system always remains in a benign state, thus guaranteeing rollback resistance. This guarantee is file system- and application-agnostic.
@@ 5. System Model
Rollbaccine maintains block device crash consistency in the presence of rollback attacks through fault-tolerant replication of disk writes. Rollbaccine consists of $N$ machines, running within TEEs. One machine is the primary where the application executes, while the remaining $N-1$ nodes are backups.
Correctness guarantees. Rollbaccine provides rollback detection up to any number of failures, and rollback resistance up to $f$ crashes, rollbacks, or other failures in line with our threat model (§ 2.2).
During execution, the primary replicates writes by broadcasting it to at least $f+1$ machines (including itself). After a crash (and potential rollback), the recovering machine’s disk is restored to a block device crash-consistent state by contacting at least $f+1$ existing machines and recovering from the most up-to-date machine. $N$ can be configured to be any value between $f+1$ and $12f+$ 1 (the largest $N$ where the replication and recovery quorums still intersect).
The $f$ failure assumption is only realistic if failures are independent. A malicious cloud provider, however, could simultaneously attack all $N$ machines. In that case, because the integrity of Rollbaccine’s recovery logic is protected by the TEE, Rollbaccine would still detect the rollback. It would simply fail to recover due to its inability to reach a quorum of existing machines. This effectively turns the rollback attack (on safety) into a denial-of-service attack (on liveness). A user could then take action against the cloud provider; knowing this, the cloud provider is incentivized against such attacks (nimble, ; secureKeeper, ; lcm, ; cyclosa, ).
Selecting $N$. The relationship between $N$ and $f$ represent a configurable tradeoff between availability and cost. Traditional consensus protocols maximize $N$ by setting it to $12f+$ 1, reducing the effect of stragglers and failures on replication and recovery at the cost of additional machines. Primary-backup and chain replication (chainReplication, ) protocols minimize $N$ by setting it to $f+1$, reducing the number of machines at the cost of availability when nodes fail.
By default, Rollbaccine sets $N=f+1$ and requires explicit recovery to recoup liveness; Rollbaccine is only live in the absence of failures. This is in line with what a user can expect from a traditional cloud deployment: if a VM crashes, the user (or some third-party software) is responsible for restarting it or deploying another VM. Rollbaccine preserves this abstraction in addition to transforming rollback attacks into benign crashes. This allows Rollbaccine to reduce the number of backups, minimizing networking bandwidth, overhead, and cost. Indeed, in our problem setting, where Rollbaccine is solely responsible for preserving the correctness of the block device, an attacker can always violate liveness by crashing the primary VM executing the user’s application (nimble, ; secureKeeper, ; lcm, ; cyclosa, ).
@@ 6. Design
Rollbaccine must balance mitigating the high cost of replication with the constraints placed on recovery by block devices. This tension manifests itself in two areas: synchronous vs asynchronous replication, and multi-threaded vs single-threaded execution.
Synchronous vs. asynchronous replication. Naïve, synchronous replication of all disk writes to a set of backups is prohibitively expensive, as it requires waiting for responses from sufficiently many backups before acknowledging each write. Asynchronous replication, on the other hand, introduces a window of vulnerability during which data may be lost: the write may have optimistically been confirmed before replicating to sufficiently many backups.
@@
Multi-threaded vs. single-threaded execution. Disks achieve high throughput by allowing writes to be processed in parallel. To maintain the multithreaded nature of disks when replicating, Rollbaccine exploits the fact that write invocations between replicas need not be processed in the same order. Because the backups’ states are only used in the event of a crash or rollback attack, they simply need to be durable cuts (Definition 5) of the primary’s state in order to achieve block device crash consistency. In other words, the backups must respect happens-before relationships and the semantics of persistence flags, but are free to reorder all other operations. Concretely, Rollbaccine backups submit write block I/Os to disk according to the total order assigned by the primary, but submissions do not block on the completion of previous I/Os unless they conflict.
@@ 6.1.1. Asynchronous writes on the primary
Writes without persistence flags, such as those before the fsync in Table 1, are asynchronously replicated to backups. When Rollbaccine intercepts a write on the primary $p$, it encrypts and hashes it (with AEAD), stores the hash in memory, then atomically (1) assigns it a monotonically increasing writeIndexp and (2) places it on the network queue, where it will be signed and sent to the backups.
Keeping integrity metadata in memory. Traditional integrity-preserving systems that keep integrity metadata on disk (dmIntegrity, ; azureStorage, ) are vulnerable to attacks that simultaneously rollback the data and its integrity metadata.
@@
Prior work has also explored using Merkle trees (without replication) to detect disk integrity violations (dm-x, ; vpfs, ; mlsdisk, ; memoir, ; ariadne, ). Their correctness rests on keeping the root/tail hash in “small trusted storage”. Even if small trusted storage were available (and evidence suggests otherwise (rote, ; nimble, )), these solutions are at best rollback detecting and not resistant; once the metadata is corrupted, it cannot be recovered.
@@
Altogether, this mechanism converts concurrent writes to the same block into sequential writes. This is similar to the approach taken in Harmonia (harmonia, ), CrossFS (crossFS, ), and dm-integrity (dmIntegrity, ), which represents the state-of-the-art in the understanding of block device semantics.
@@ 6.1.2. Asynchronous writes arriving at the backups
The naïve solution, executing all writes one-after-the-other according to writeIndex, is a non-starter performance-wise. The challenge is then parallelizing these writes safely. To do so, the backups need to determine which writes are to the same block, as block semantics allows non-conflicting writes to be ordered arbitrarily (crashmonkey, ; crashMonkeyAndACE, ; dmIntegrityMailingList, ).
@@
In order to preserve happens-before relationships between writes to the same block, the backups must still submit writes to disk in order of writeIndexp as assigned by the primary, but do not wait for the disk to finish processing previous writes; only conflicting writes need to block. Concretely, once a backup $b$ receives a write with $\texttt{writeIndex}_{p}=\texttt{writeIndex}_{b}+1$, it atomically increments writeIndexb, then follows the same process depicted in Figure 3. This simultaneously allows disk bandwidth to be fully utilized on the backup, allowing non-conflicting writes to be concurrently in-flight, while preserving write ordering over individual blocks.
@@ 6.1.3. Synchronous writes
Writes tagged with persistence flags are handled identically with one exception: Rollbaccine does not return the write until backups confirm that they have received all writes with a lower write_index.
@@ 6.1.4. Reads
Reads are performed on the primary and do not involve the backups. To maintain integrity for concurrent reads and writes to the same block, Rollbaccine inserts reads in the same pending queue/involved tree as concurrent writes. Once the read can be executed, Rollbaccine fetches the corresponding page from disk, decrypts it, checks it against the hash in memory, and returns the decrypted page if the integrity check succeeds. If the check fails—because of a rollback attack or a benign disk corruption—Rollbaccine crashes the machine, entering recovery upon restarting, where the backups are used to verify the integrity of the primary’s disk and recover corrupted data.
@@ 6.2. Recovery
Rollbaccine’s recovery protocol differs from traditional disk recovery in two ways. First, it must retrieve in-memory integrity metadata and any corrupted disk pages from the most up-to-date backup. Second, it must prevent split-brain attacks, where an attacker could feign a crash, wait for the user to “restart” the “crashed” machine while actually starting a new machine, then route external client traffic between the new and “crashed” machines as desired (narrator, ).
We prevent split-brain attacks during recovery by drawing an equivalence to reconfiguration (verticalPaxos, ; paxosComplex, ). We require the client (or some fault-tolerant third party) to provide each restarted machine a new identity, even if the physical hardware is the same. Each recovering machine then joins a new configuration that excludes its crashed self, ensuring that stale machines no longer participate in the protocol.
Rollbaccine’s recovery protocol is based on Matchmaker Paxos (matchmakerPaxos, ), a state-of-the-art vertical reconfiguration (verticalPaxos, ) protocol that uses two round-trips: one to a fault-tolerant third party to establish the current configuration (the active primary and backups), and another to invalidate all previous configurations. We use CCF (ccf, ), a TEE-based distributed key-value store, as the third party. Relying on a third party for reconfiguration is a practice common among consensus protocols (verticalPaxos, ; chainReplication, ; boxwood, ; gfs, ; pacifica, ) and does not affect the critical path of Rollbaccine. As CCF is only utilized during recovery, a single CCF service could support many Rollbaccine instances.
During recovery, the recovering node identifies the most up-to-date node—the designated node with the latest configuration and highest writeIndex—then copies the designated node’s integrity metadata, scans its local disk, and copies over any corrupted pages. The recovering node then alerts any nodes in its new configuration, which copy their integrity metadata and corrupted pages from the recovering node as well in order to maintain consistency. Once the disk is repaired, Rollbaccine is mounted and can be used as-is.
Details of the recovery protocol are in Appendix A.

### v2

@@ 4. Block Device Crash Consistency
Rollbaccine provides rollback resistance by guaranteeing that, following a rollback attack, the disk always recovers to a safe state, one that it could have recovered to after a benign crash. A formal definition of safe state, independent of file system semantics, is thus necessary. Concretely, we formalize block device consistency as the states to which a block device (i.e. a disk) can recover to after a crash.
We begin with the assumption that disk read and write operations ($O$) are atomic, which is consistent with prior work and the inherent properties of disks (brown2002oracle; sivathanu2003semantically; pillai2014all). We also assume that writes to block devices are guaranteed to persist across crashes only when the persistence flags REQ_FUA or REQ_PREFLUSH are used (crashmonkey; crashMonkeyAndACE; vijayThesis; fuaPreflush). REQ_FUA guarantees that when the write is completed, it must be persisted, whereas REQ_PREFLUSH guarantees that any previously completed operation is persisted.
@@
Read requests to block $b$ by thread $t$ are written $R_{inv,t}(b)$; responses are $R_{res,t}(b,val)$, where valval is the value returned. Write requests are $W_{inv,t}(b,val,sync)$, where the value to write (if any) is val and sync is a tag with one of the following values: REQ_FUA, REQ_PREFLUSH, REQ_FUA|REQ_PREFLUSH, or $\varnothing$. $W_{res,t}(b)$ is the matching response. We assume that blocks are always written to before they are read from.
To define block device crash consistency, we take as our starting point the definitions of Izraelevitz et al (durableLinearizability). We will build up to a definition of linearizability before extending it to crashes. We start by defining a sequential history.
@@
This allows us to reason about multi-threaded histories by comparing each thread’s execution to a sequential history. [^2] For programs that issue concurrent operations per thread using async I/O, we can map each physical thread to multiple abstract threads. When multiple threads operate over the same block, we use the happens-before relationship to order writes and reads on different threads, as this is necessary to determine whether a history satisfies reads-see-writes.
@@
(3) $V_{1}=O_{res}(b)$ and $V_{2}=W_{inv}(b^{\prime},val,sync)$ where syncsync contains REQ_PREFLUSH, or
@@
Criteria 1 and 4 are standard (linearizability); Criterion 2 states that crashes are global events; all events either happen-before or after a crash (durableLinearizability). Criterion 3 is new and captures the global semantics of REQ_PREFLUSH: once a REQ_PREFLUSH is invoked, the operation only completes when all previous writes from every thread are flushed and persisted.
The happens-before relationship allows us to define what each read returns. Each read of block $b$ must return the value of the latest completed write to that same block $b$, as long as there are no crashes in-between (during which writes may be lost). We formalize this as the reads-see-writes property, which is only defined for crash-free periods of history (we call these eras $\mathcal{E}$).
@@
A history $\mathcal{H}$ respects reads-see-writes if $\forall R_{res}(b,val)\in\mathcal{H}$, there is a preceding write invocation $W_{inv}(b,val,sync)$ with that same valval such that $\mathcal{H}=\mathcal{H}_{0}W_{inv}\mathcal{E}_{0}W_{res}\mathcal{E}_{1}R_{inv}\mathcal{E}_{2}R_{res}\mathcal{H}_{1}$, and there does not exist another $W_{inv}(b)$ in the eras $\mathcal{E}_{0},\mathcal{E}_{1},\mathcal{E}_{2}$.
Finally, we consider pending invocations: invocations without a matching response. Pending writes in particular require care as they may (or may not) have been processed by the underlying block device and reflected in the next read. We write compl()($\mathcal{H}$) to be the set of histories generated from $\mathcal{H}$ by inserting matching responses after some pending invocations. This models situations where pending operations have been persisted to disk. In contrast, let trunc()($\mathcal{H}$) be the history generated from $\mathcal{H}$ by removing all pending invocations. This reflects histories where the operation was not persisted.
@@
(1) $\mathcal{S}$ respects reads-see-writes
(2) $\forall t$, $\mathcal{H}^{\prime}[t]=\mathcal{S}[t]$ (i.e. $\mathcal{H}^{\prime}$ and $\mathcal{S}$ are equivalent)
(3) $V_{1}\prec V_{2}$ in $\mathcal{H}^{\prime}$ implies $V_{1}\prec V_{2}$ in $\mathcal{S}$.
@@
(1) if $\mathcal{H}^{\prime}$ contains $W_{inv}(b,val,sync)$ and its matching response $W_{res}(b)$ where syncsync contains REQ_FUA or REQ_PREFLUSH, then $\mathcal{D}$ must contain $W_{res}(b)$,
@@
The durable cut is a cut of history that contains (1) all writes tagged with persistence flags and (2) any writes that happen-before a write already in the cut. This is where the extra criteria for REQ_PREFLUSH in the happens-before relationship becomes relevant; if a REQ_PREFLUSH completed, then it must be in the durable cut, and any operations that happened-before it must be in the cut as well.
@@
Assuming only benign system crashes and no random disk corruption, block device crash consistency precisely captures the set of histories produced by a disk (dmIntegrityMailingList). We prove that all histories produced by Rollbaccine are block device crash consistent in Appendix B, ensuring that the system always remains in a safe state and is thus rollback resistant. This guarantee is file system and application-agnostic.
@@ 5. System Model
Rollbaccine maintains block device crash consistency in the presence of rollback attacks through fault-tolerant replication of disk writes.
Rollbaccine consists of $N$ machines, running within TEEs. One machine is the primary where the application executes, while the remaining N1N-1 nodes are backups. During execution, the primary replicates writes to at least $f$ backups. After a crash (and potential rollback), the recovering machine’s disk is restored to a block device crash-consistent state by contacting at least NfN-f existing machines and recovering from the most up-to-date machine.
Selecting $N$. NN can be configured to be any value between $f+1$ and $2f+1$ (the largest $N$ where the replication and recovery quorums still intersect). The relationship between $N$ and $f$ represent a configurable tradeoff between availability and cost. Traditional consensus protocols maximize availability by setting $N=2f+1$, reducing the effect of failures on recovery (remaining live with up to $f$ failures) at the cost of additional machines. Primary-backup and chain replication (chainReplication) protocols minimize cost by setting $N=f+1$, reducing the number of machines at the cost of availability when nodes fail.
By default, Rollbaccine sets $N=f+1$ and requires explicit recovery to recoup liveness; Rollbaccine is only live in the absence of failures. This is in line with what a user can expect from a traditional cloud deployment: if a VM crashes, the user (or some third-party software) is responsible for restarting it or deploying another VM. Setting $N=2f+1$ guarantees Rollbaccine’s liveness with up to $f$ backup failures.
@@ 6. Design
Rollbaccine must balance the high cost of replication with the constraints placed by block devices. This tension manifests itself in two areas: synchronous vs asynchronous replication, and multi-threaded vs single-threaded execution.
Avoiding unnecessary synchronous replication. Naïve, synchronous replication of all disk writes to a set of backups is prohibitively expensive, as it requires waiting for responses from sufficiently many backups before acknowledging each write. Asynchronous replication, on the other hand, introduces a window of vulnerability during which data may be lost: the write may have optimistically been confirmed before replicating to sufficiently many backups.
@@
Multi-threaded execution for non-conflicting operations. Disks achieve high throughput by allowing writes to be processed in parallel. To maintain the multithreaded nature of disks when replicating, Rollbaccine exploits the fact that write invocations between replicas need not be processed in the same order. Because the backups’ states are only used in the event of a crash or rollback attack, they simply need to be durable cuts (Definition 5) of the primary’s state in order to achieve block device crash consistency. In other words, the backups must respect happens-before relationships and the semantics of persistence flags, but are free to reorder all other operations. Concretely, Rollbaccine backups submit write block I/Os to disk according to the total order assigned by the primary, but submissions do not block on the completion of previous I/Os unless they conflict.
@@ 6.1.1. Asynchronous writes on the primary
Writes without persistence flags, such as those before the fsync in Table 1, are asynchronously replicated to backups. When Rollbaccine intercepts a write on the primary $p$, it encrypts and hashes it (with authenticated encryption), stores the hash in memory, then atomically (1) assigns it a monotonically increasing writeIndexp and (2) places it on the network queue, where it will be signed and sent to the backups.
Keeping integrity metadata in memory. Traditional integrity-preserving systems that keep integrity metadata on disk (dmIntegrity; azureStorage) are vulnerable to attacks that simultaneously rollback the data and its integrity metadata.
@@
Prior work has also explored using Merkle trees (without replication) to detect disk integrity violations (dm-x; vpfs; mlsdisk; memoir; ariadne; fastver). Their correctness rests on keeping the root/tail hash in “small trusted storage”. Even if small trusted storage were available (and evidence suggests otherwise (rote; nimble)), these solutions are at best rollback detecting and not resistant; once the metadata is corrupted, it cannot be recovered.
@@
Altogether, this mechanism converts concurrent writes to the same block into sequential writes. This is similar to the approach taken in Harmonia (harmonia), CrossFS (crossFS), and dm-integrity (dmIntegrity), which represents the state-of-the-art in the understanding of block device semantics.
@@ 6.1.2. Asynchronous writes arriving at the backups
The naïve solution, executing all writes one-after-the-other according to writeIndex, is a non-starter performance-wise. The challenge is then parallelizing these writes safely. To do so, the backups need to determine which writes are to the same block, as block semantics allows non-conflicting writes to be ordered arbitrarily (crashmonkey; crashMonkeyAndACE; dmIntegrityMailingList).
@@
In order to preserve happens-before relationships between writes to the same block, the backups must still submit writes to disk in order of writeIndexp as assigned by the primary, but do not wait for the disk to finish processing previous writes; only conflicting writes need to block. Concretely, once a backup $b$ receives a write with $\texttt{writeIndex}_{p}=\texttt{writeIndex}_{b}+1$, it atomically increments writeIndexb, then follows the same process depicted in Figure 3. This simultaneously allows non-conflicting writes to be concurrently in-flight while preserving write ordering over individual blocks.
@@ 6.1.3. Synchronous writes
Writes tagged with persistence flags are handled identically with one exception: Rollbaccine does not return the write until backups confirm that they have received all writes with a lower writeIndex.
@@ 6.1.4. Reads
Reads are performed on the primary and do not involve the backups. To maintain integrity for concurrent reads and writes to the same block, Rollbaccine inserts reads in the same pending queue/involved tree as concurrent writes. Once the read can be executed, Rollbaccine fetches the corresponding page from disk, decrypts it, checks it against the hash in memory, and returns the decrypted page if the integrity check succeeds. If the check fails—because of a rollback attack or a benign disk corruption—Rollbaccine crashes the machine, entering recovery upon restarting.
@@ 6.2. Recovery
Rollbaccine’s recovery protocol differs from traditional disk recovery in two ways. First, it must retrieve in-memory integrity metadata and any corrupted disk pages from the most up-to-date backup. Second, it must prevent split-brain attacks, where an attacker could feign a crash, wait for the user to “restart” the “crashed” machine while actually starting a new machine, then route external client traffic between the new and “crashed” machines as desired (narrator).
We prevent split-brain attacks during recovery by drawing an equivalence to reconfiguration (verticalPaxos; paxosComplex). We require the client (or some fault-tolerant third party) to provide each restarted machine a new identity, even if the physical hardware is the same. Each recovering machine then joins a new configuration that excludes its crashed self, ensuring that stale machines no longer participate in the protocol.
Rollbaccine’s recovery protocol is based on Matchmaker Paxos (matchmakerPaxos), a state-of-the-art vertical reconfiguration (verticalPaxos) protocol that uses two round-trips: one to a fault-tolerant consensus protocol to establish the current configuration (the active primary and backups), and another to invalidate all previous configurations.
We use CCF (ccf) as the consensus protocol as opposed to implementing our own. CCF is a mature, fault-tolerant, rollback-resistant, replicated key-value store that requires no additional trust assumptions. As CCF is only needed at the beginning of recovery, its performance does not affect Rollbaccine; it can be configured to run on the same machines as Rollbaccine or shared across many Rollbaccine instances to minimize cost.
During recovery, the recovering node contacts at least NfN-f nodes from the latest configuration, identifies the node with the highest writeIndex, copies that node’s integrity metadata, scans its local disk, and copies over any corrupted pages. The recovering node then alerts any nodes in its new configuration, which copy their integrity metadata and corrupted pages from the recovering node as well in order to maintain consistency. Once the disk is repaired, Rollbaccine is mounted and can be used as-is.
If fewer than NfN-f nodes from the latest configuration can be reached during recovery, then Rollbaccine assumes the worst—that its disk has been rolled-back by a Type II attacker—and aborts recovery.
The full reconfiguration protocol and its proof are in Appendix A.